The probability of snow is 0.03. What is the probability of no snow?
Answer by jim_thompson5910(35256)

P(Snow) = 0.03

P(No Snow) = 1 - P(Snow) ... since P(Snow) + P(No Snow) = 1

P(No Snow) = 1 - 0.03

P(No Snow) = 0.97


So the probability of no snow is 0.97